OpenStudy (anonymous):

Write the equation of the circle in standard form. Find the center, radius, intercepts. x^2 + y^2 - 8x - 6y + 9= 0

OpenStudy (bahrom7893):

just put this into the form (x-h)^2+(y-k)^2 = r^2.. soo if we do some arithmetic/algebra..

OpenStudy (bahrom7893):

First get the xs and ys together into their groups x^2-8x + y^2-6y + 9 = 0

OpenStudy (bahrom7893):

(x^2 - 2 * 4 * x) + (y^2 - 2 * 3 * y) + 9 = 0

OpenStudy (anonymous):

math makes me want to die

OpenStudy (bahrom7893):

Complete the square: (x^2 - 2 * 4 * x + 4^2) + (y^2 - 2 * 3 * y + 3^2) + 9 - 4^2 - 3^2 = 0

OpenStudy (anonymous):

does * = multiply?

OpenStudy (bahrom7893):

Yea (x-4)^2 + (y-3)^2 + 9 - 16 - 9 = 0 (x-4)^2 + (y-3)^2 - 16 = 0 (x-4)^2 + (y-3)^2 = 16

OpenStudy (bahrom7893):

which can be rewritten as: (x-4)^2 + (y-3)^2 = 4^2

OpenStudy (anonymous):

im just going to start emailing you my homework. end of discussion.

OpenStudy (bahrom7893):

So that's in the standard form: (x-h)^2 + (y-k)^2 = r^2

OpenStudy (bahrom7893):

hey that's not very nice!

OpenStudy (bahrom7893):

The center is (h;k) and the radius is r

OpenStudy (anonymous):

I would say Hi too! :D like..." Hi! Love ya....can I have answers by 2am EST? thankkksss!"\

OpenStudy (bahrom7893):

So final ans: Circle in standard form: (x-4)^2 + (y-3)^2 = 4^2 Center: (4;3) Radius: 4

OpenStudy (bahrom7893):

i'd be asleep, so i would only read the email at 7am.. and respond at around 10pm

OpenStudy (anonymous):

Im going to be up for a while reading this in my book and making sense of it. I take math online so I don't have much help

OpenStudy (anonymous):

Its ok Bah Im not really going to email my homework lol.

OpenStudy (bahrom7893):

ik.. just tellin you what would happen if you did lo

OpenStudy (anonymous):

thanks for the help! <3